OpenStudy (anonymous):

Algebra Questions: 3. Simplify the radical expression. square root 150x^3k^4 A) 75xk^2 B) 5xk square root 6xk C) 5xk^2 square root 6x D) 5 square root 6x^3k^4

OpenStudy (anonymous):

@SkepticGod

OpenStudy (anonymous):

Break 150 down into prime numbers and squares. If it were the number 98: 2*49=98 49 is 7^2, so you can pull 7 out of the radical. Then just pull each variable to the second power out.

jhonyy9 (jhonyy9):

150 = 25*6 = 5^2*6 x^3 = x^2*x k^4 = k^2*k^2 - so do you see the perfect squares - take them out from squaroot and check what will get

jhonyy9 (jhonyy9):

@Answers101 you have said me that you understand it so than wait

jhonyy9 (jhonyy9):

please you cooperate for i can help you

jhonyy9 (jhonyy9):

understanding it sure

jhonyy9 (jhonyy9):

sqrt150 = sqrt(25*6)= sqrt(5^2*6) = 5sqrt6 yes ?

jhonyy9 (jhonyy9):

continue it please